# Trust, But Verify: A Survey of Randomized Smoothing Techniques

### v1

@@ Abstract
Machine learning models have demonstrated remarkable success across diverse domains but remain vulnerable to adversarial attacks. Empirical defence mechanisms often fall short, as new attacks constantly emerge, rendering existing defences obsolete. A paradigm shift from empirical defences to certification-based defences has been observed in response. Randomized smoothing has emerged as a promising technique among notable advancements. This study reviews the theoretical foundations, empirical effectiveness, and applications of randomized smoothing in verifying machine learning classifiers. We provide an in-depth exploration of the fundamental concepts underlying randomized smoothing, highlighting its theoretical guarantees in certifying robustness against adversarial perturbations. Additionally, we discuss the challenges of existing methodologies and offer insightful perspectives on potential solutions. This paper is novel in its attempt to systemise the existing knowledge in the context of randomized smoothing.
Index Terms:
Deep learning, adversarial attacks, randomized smoothing, certification-based defense
publicationid: pubid: *Equal Contribution
I Introduction
@@
Machine Learning (ML) has significantly advanced, notably since the development of deep neural networks (DNNs) [79]. However, a significant challenge is the susceptibility of state-of-the-art networks to adversarial examples [65, 68] limiting their application in safety-critical areas [77]. Adversarial attacks include techniques like DeepFool [95], AutoAttack[96], and patch-based attacks[97]. Existing defenses against adversarial attacks have proven ineffective against stronger methods, creating a continual cat-and-mouse game between attackers and defenders where attackers continually devise stronger attacks [81], and defenders attempt to provide robustness against these attacks citeb111. Empirical and heuristic defenses include denoising autoencoders [98], using training-based defenses like adversarial training [99], and defensive distillation[100]. Despite efforts to fortify ML models, it has been suggested that creating DNNs inherently robust to adversarial examples may be inevitable [101], [102]. Some approaches include those explicitly based on the architecture of the DNN or otherwise [106, 84]. Further, it has been shown that enhancing robustness might come at the cost of a drop in accuracy [73, 78]. This ongoing struggle has led researchers to explore certified robustness as an alternative.
I-A Adversarial Examples
Adversarial examples are subtle input perturbations causing the classifier to give incorrect outputs. Consider a classifier from $\mathbb{R}^{d}$ to classes $C$, mathematically defined as, $f:\mathbb{R}^{d}\rightarrow[C]$, that assigns a label to an input $x\in X$, we define an adversarial example $x_{0}$ as one where $D(x,x_{0})<\epsilon$ for a small $\epsilon>0$, and $f(x)\neq f(x_{0})$. Here, $D(\cdot,\cdot)$ represents a distance metric, often expressed using an $l_{p}$ norm, with $l_{1}$, $l_{2}$ and $l_{\infty}$ norms being the most commonly used in the literature.
I-B Empirical and Certified Robustness
Given a classifier $f$, We say $f$ is adversarially robust if, when provided with an input $x$ with the actual label $y$, it can consistently classify all inputs within a certain $p$-norm ball ($B_{p,r}(x)$) of radius $r$ centered at $x$ correctly. In simpler terms, $f$ is robust if it correctly classifies nearby inputs. Further, we define empirical and certified defenses. Empirical defenses against adversarial examples refer to strategies implemented to safeguard machine learning models from adversarial examples. These empirical strategies are developed and validated based on experimental findings rather than theoretical guarantees. In contrast, certified defenses against adversarial examples refer to strategies that provide mathematically provable assurances that a machine learning model can withstand adversarial attacks within certain constraints. Empirical robustness is often determined by attempting to find inputs ($x_{0}$) within the ball $B_{p,r}(x)$ that are misclassified by $f$, providing an upper estimate of the classifier’s actual robustness. On the other hand, certified robustness serves as a provable lower bound on the classifier’s robustness, ensuring a certain level of correctness within the specified region.
I-C Randomized Smoothing
@@
Randomized smoothing [72, 2] has emerged as a powerful certification-based defense against adversarial attacks. The fundamental idea behind randomized smoothing is to create a smoothed classifier by applying a convolution with Gaussian noise to the base classifier. Given a base classifier $f$, We create a smoothed version of this classifier, denoted as $g$, which predicts the most likely class $c$ that the base classifier $f$ will predict for a noised version of the input $x+\epsilon$, where $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$. Thus, $g$ can be defined as:
$g(x+\epsilon)=\arg\max_{c\in[C]}P(f(x+\epsilon)=c)\text{, where }\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$ (1)
Figure 1: This figure showcases the decision boundaries for $l_{1}$, $l_{2}$, and $l_{\infty}$ norms, visually representing their sensitivity to adversarial perturbations. The diamond, circle, and square-shaped boundaries signify $l_{1}$, $l_{2}$, and $l_{\infty}$ norms, respectively, showcasing the technique’s impact on different normative spaces for a given certified radius.
Lecuyer et al. [72] and Li et al. [2] showed that the smoothed classifier $g$ will consistently classify within a certified radius around the input $x$ under $l_{2}$ norm considerations. However, the guarantees provided by these papers were loose due to a noticeable gap between the theoretical lower bounds and the empirical accuracy, suggesting the existence of a tighter radius around the input $x$ than those predicted.
Cohen et al. [1] was the first paper to prove tight robustness guarantees for a randomized smoothing classifier against adversarial attacks constrained under the $l_{2}$ norm. Given the top two classes, $c_{A}$ and $c_{B}$, that the base classifier is most likely to predict for the sample, the robust radius $r$ is given by:
$r=\frac{\sigma}{2}(\Phi^{-1}(p_{A})-\Phi^{-1}(p_{B}))$ (2)
where $p_{A}=P(f(x+\epsilon)=c_{A})$ and $p_{B}=P(f(x+\epsilon)=c_{B})$, with the condition that $p_{A}>0.5$ and $\Phi$ represents the CDF of the standard Gaussian distribution. This demonstrates that if $g$ correctly predicts the input $x$, we can provide provable $l_{2}$ robustness of our model in a probabilistic sense. However, Cohen et al. [1] highlighted that it is practically not possible to calculate the exact values of the top two class probabilities, i.e., $p_{A}$ and $p_{B}$. Hence, they introduce a modified version of the above statement that utilizes the Clopper-Pearson bounds for $p_{A}$ and $p_{B}$. For $p_{A}$, the Clopper-Pearson lower bound, denoted as $\underline{p_{A}}$, and for $p_{B}$, the Clopper-Pearson upper bound, denoted as $\overline{p_{B}}$, can be calculated using Monte-Carlo sampling. To theoretically prove that using this combination of Clopper-Pearson and Monte-Carlo sampling will give us a lower bound on the actual certifiable radius $r$, they use Neyman-Pearson lemma (NP lemma). Furthermore, to reduce time complexity, they approximate $\overline{p_{B}}$ with $1-\underline{p_{A}}$, meaning that if the base classifier correctly classifies the input $x$, then any example $x_{0}$ within the $l_{2}$ distance of x, $D(x,x_{0})<\sigma\Phi^{-1}(\underline{p_{A}})$ will be correctly classified by the smoothed classifier $g$.
Following Cohen et al.’s seminal work [1], the concept of randomized smoothing has spurred a proliferation of studies. This paper endeavours to provide a comprehensive summary of these works, elucidating significant challenges, notable advancements and extracting key insights. The organization of the survey is structured into three main sections. Section II critically examines the major issues associated with randomized smoothing, categorizing these challenges into three broad domains. Subsequently, a detailed discussion ensues on these challenges, followed by an exploration of future research directions.
Moving to Section III, we delve into significant advancements that focus on enhancing radius certifications beyond those outlined in Cohen et al. [1]. This section also addresses the challenges presented in Section II, providing discussions for the same. Section IV broadens the perspective by discussing applications of randomized smoothing, which extend beyond the original focus of Cohen et al. [1], which was primarily limited to image classification tasks.
@@
Rethinking Scalability in Randomized Smoothing: An Analytical Perspective Due to its theoretical guarantees, randomized smoothing is commonly hailed as a crucial certification-based defense mechanism. However, through our analysis, we raise a fundamental question: Despite all the theoretical assurances, is randomized smoothing truly scalable? Throughout our examination, we highlight the limitations of existing works that predominantly emphasize theoretical guarantees while overlooking the critical aspect of computational complexity, especially during inference, and dimensionality considerations. This concern holds significant weight, particularly for machine learning applications requiring minimal inference time, where randomized smoothing may fall short as a practical solution. Hence, in addition to serving as a survey of randomized smoothing techniques, we scrutinize each work from the perspective of scalability. At the end of each section, we offer insights and discussions on how the various works discussed within that section may influence the scalability of randomized smoothing.
II Challenges
Randomized smoothing is a powerful architecture-independent technique for certified robustness, and it has only gained popularity over the years; however, as with any other method, it also has its limitations. In the following subsections we define and discuss three significant challenges in RS that commonly encountered in literature. These challenges affect the scalability and applicability of RS in many domains, which has currently offered promising empirical results in certified robustness.
II-A Curse of Dimensionality
@@
In their work, Cohen et al. [1] left the case of general $l_{p}$ norms as an open problem, suspecting that smoothing with other noise distributions might provide robustness guarantees for these cases, and only discussed randomized smoothing in the context of $l_{2}$ norms. Additionally, they state that, despite the certified radius being independent of the input dimension $d$, randomized smoothing can be easily scaled to higher dimension images as in the case of higher dimension images, we can increase the variance of the noise, which will eventually lead to increase in the robust radius. This approach has its issues, as will be discussed in the next section. Later, two independent works, Blum, Avrim, et al. [7] and Kumar, Aounon, et al. [11], showed that as the dimension of the input increases, it becomes increasingly difficult to defend against the general $l_{p}$ norm cases for $p>2$ as the certified radius decreases.
Blum, Avrim, et al. [7] showed that, given any noise distribution $\mathcal{D}$ and the noise vector $\epsilon\sim\mathcal{D}$, 99% pixels of the noise vector, must satisfy $\mathbb{E}\epsilon_{i}^{2}=\Omega(d^{1-\frac{2}{p}}r^{2}\frac{1-\delta}{\delta^{2}})$, where $r$ is the robust radius and $\delta$ is the difference between the probability scores of the top two classes predicted by the smoothed classifier (highest score-class $c_{A}$ and runner-up class $c_{B}$). They further proved that the problem of finding the certified radius in the $l_{\infty}$ case can be approximated as finding the $l_{2}$ radius times $1/\sqrt{d}$ of that radius.
Subsequently, Kumar, Aounon, et al. [11] show, in the specific case of a generalized Gaussian distribution, tighter bounds than those attained by Blum, Avrim, et al. [7]. Their work states that the certified radius for an $l_{p}$ norm decreases as $O\left(d^{\frac{1}{2}-\frac{1}{p}}\right)$ with data dimension $d$ for $p>2$. The family of distributions discussed in this paper includes the Laplacian, Gaussian, and uniform distributions, commonly employed in literature for randomized smoothing techniques. Wu, Yihan, et al. [75] then extend the discussion for the case of spherical symmetric distributions (which includes the Gaussian distribution) and present upper bounds for $p>2$. It also discusses upper bounds for $p=2$ for many popular smoothing distributions, which have yet to be discussed. Yang, Greg, et al. [4] go into more detail about $l_{1}$, $l_{2}$, and $l_{\infty}$ threat models to derive from the class of i.i.d. distributions, the most optimal smoothing distribution using the concept of Wulff Crystals. Overall, their results in this regard can be summarised as follows: the uniform distribution is the most optimal choice for $l_{1}$ attacks, the Gaussian distribution, as mentioned by Cohen et al. [1], is the most optimal for $l_{2}$ attacks, and that for $l_{\infty}$, the Gaussian distribution performs best using appropriate approximations. There is a tight $O\left(\min\left(1,d^{\frac{1}{2}-\frac{1}{p}}\right)\right)$ dimensionality bound for $l_{p}$ norms without using any additional information other than the Neyman-Pearson technique.
Ettedgui, Raphael, et al. [66] provide a different perspective and argue about the information-theoretic limitations faced by RS and how they are not intrinsic but a byproduct of the current certification methods. They further discuss that these certificates need to include more information about the classifier and ignore the local curvature of the decision boundary, which leads to sub-optimal robustness guarantees as the dimension of the problem increases. It is possible to overcome this by generalizing the Neyman-Pearson Lemma and collecting more information about the classifier. The paper shows that it is possible to approximate the optimal certificate with arbitrary precision by probing the decision boundary with several noise distributions. More specifically, the paper achieves this without sampling in high dimensions by combining uniform and Gaussian distribution and leveraging the isotropic properties of the latter. Further, this process retains natural accuracy as it is executed at certification time. The paper then gives theoretical insight into how to mitigate the computational cost of a classifier-specific certification.
Pfrommer, Samuel et al. [41] present notable work in the sense that they implicitly propose a solution to the curse of dimensionality, suggested by Blum, Avrim, et al. [7]. It combines the limitations of randomized smoothing in light of the curse of dimensionality with the manifold hypothesis [74] and shows that using various dimensionality reduction techniques improves robust guarantees in theory, as opposed to heuristic approaches in previous works by Mustafa, Aamir, et al. [76] and Kurakin, Alexey et al. [77].
@@
So far, only Li, Linyi, et al. [3] have provided proofs and a possible solution to the issue of the curse of dimensionality using a technique called double sampling randomized smoothing (DSRS). They utilize a primary distribution in the form of a general Gaussian distribution and additional information in the form of a sampled probability from a secondary truncated distribution that is not necessarily different from the primary distribution. For $l_{2}$ norm, they provide a tighter radius and prove that their method can certify $\Theta\left(\sqrt{d}\right)$ robust radius bounds. As mentioned by Cohen et al. [1], for the $l_{\infty}$ case, the certified radius is equivalent to $1/\sqrt{d}$ times the $l_{2}$ robust radius, which on using [3] would remove the dependency on the dimensions of the input.
We note that there has been an in-depth theoretical exploration of the issue termed the curse of dimensionality. The base conclusion of all papers remains that there is a tight bound on certified radii that decreases with increasing dimensionality, which arises from statistical, probabilistic, and information-theoretic perspectives based on the existing methodology for RS. While we certainly encourage more research into that aspect, we notice the dire lack of solutions for the same. Apart from DSRS, no major breakthrough has been achieved to solve this issue. As a result, we encourage works to explore various solutions to this issue.
II-B Robustness-Accuracy Trade-off
@@
In adversarial robustness literature, not limited to RS, there is a focus on increasing the robustness bounds of a classifier. While this seems to be an optimal goal, there has been literature suggesting that increasing robustness comes with a tradeoff in accuracy [73], [78]. It has been studied for adversarial robustness in much detail, including theoretical considerations. In [78], the authors study the accuracy costs of improving robustness in a model. This was further explored in [73] and [80]. Both empirically and theoretically, it is seen that robustness and accuracy exist in a trade-off, contrary to what one would expect intuitively. Randomized smoothing is also vulnerable to this trade-off, and there is much to explore in this regard, leaving this an important open aspect of research for randomized smoothing techniques. Naturally, this can be seen in RS as well. Randomized smoothing involves convoluting, or smoothing, the base classifier with Gaussian noise. While this provides certified robustness, it can be inferred that there is some loss in accuracy due to noisy inputs. Termed as the robustness-accuracy tradeoff, we explore this issue concerning RS in this subsection.
Specific to RS, the theoretical and numerical aspects of Gaussian smoothing are explored in the work by Mohapatra, Jeet, et al. [51], along with which they provide theoretical results showing the limitations of RS in terms of classification accuracy. Mainly, they characterize and identify the conditions under which Gaussian smoothing leads to a decrease in classification accuracy, provide theoretical lower bounds for the magnitude of this effect, numerically inspect the behaviour of the certified radius, and use tools from information theory to analyse the effects of Gaussian smoothing during training augmentation, concluding that it leads to information loss and finally validate their results empirically.
A theoretical analysis of robustness-accuracy in terms of the benign risk associated with applying randomized smoothing on a classifier trained using noise augmentation for the base classifier is given by Pal, Ambar et al. [33]. This arises from the observation of improved performance when using a noise-augmented base classifier instead of one without any such augmentation. They derive an upper bound for any data distribution used for smoothing, which suggests randomized smoothing harms the classifier’s accuracy. This practice has pervaded most literature with ample empirical support. The paper then suggests certain distributions where smoothing may be beneficial based on the separation regime. Also, it suggests optimisation for randomized smoothing, using different noise parameters to train the base classifier. They also discuss the theoretical intuition behind this, viz., the empirical observation that real data lies on a lower dimension than the actual data dimension.

### v2

# Rethinking Randomized Smoothing from the Perspective of Scalability

@@ Abstract
Machine learning models have demonstrated remarkable success across diverse domains but remain vulnerable to adversarial attacks. Empirical defense mechanisms often fail, as new attacks constantly emerge, rendering existing defenses obsolete, shifting the focus to certification-based defenses. Randomized smoothing has emerged as a promising technique among notable advancements. This study reviews the theoretical foundations and empirical effectiveness of randomized smoothing and its derivatives in verifying machine learning classifiers from a perspective of scalability. We provide an in-depth exploration of the fundamental concepts underlying randomized smoothing, highlighting its theoretical guarantees in certifying robustness against adversarial perturbations and discuss the challenges of existing methodologies.
1 Introduction
@@
Machine learning (ML) has advanced significantly, notably since the development of deep neural networks (DNNs) ([27]). However, a significant challenge is the susceptibility of state-of-the-art networks to adversarial examples, which limits their application in areas of critical safety ([54, 15, 25]). Adversarial attacks include techniques such as DeepFool, AutoAttack, and patch-based attacks ([42, 9, 5]). Existing defenses against adversarial attacks have proven ineffective against stronger methods, creating a constant cat-and-mouse game between attackers and defenders where attackers continually devise stronger attacks and defenders attempt to provide robustness against these attacks ([6, 36]). Empirical and heuristic defenses include denoising auto-encoders, using training-based defenses such as adversarial training, and defensive distillation ([16, 38, 46]). This ongoing struggle has led researchers to explore certified robustness as an alternative. In the field of certified defense, randomized smoothing has emerged as a powerful certification-based defense against adversarial attacks.
1.1 Randomized Smoothing
@@
Randomized smoothing was introduced as a certification-based defense technique by [8]. The fundamental idea behind randomized smoothing is to create a smoothed classifier by applying a convolution with Gaussian noise to the base classifier. Given a base classifier $f$, we create a smoothed version of this classifier, denoted as $g$, which predicts the most likely class $c$ that the base classifier $f$ will predict for a noisy version of the input $x+\epsilon$, where $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$. Thus, $g$ can be calculated as:
$g(x)=\arg\max_{c\in[C]}P(f(x+\epsilon)=c)\text{, where }\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$ (1)
While [28] and [33] showed that the smoothed classifier $g$ will consistently classify within a certified radius around the input $x$ under $l_{2}$ norm considerations, [8] was the first paper to demonstrate tight robustness guarantees for a randomized smoothing classifier against adversarial attacks constrained under the norm $l_{2}$. For empirical calculation they made use of Clopper-Pearson bounds based on the Neyman-Pearson Lemma by making use of Monte Carlo Sampling.
Since Cohen et al.’s seminal work, there has been a large number of works which have been published with regards to randomized smoothing. Due to its theoretical guarantees, randomized smoothing is commonly hailed as a crucial certification-based defense mechanism. However, in our work through analysis, we raise a fundamental question:
Despite all the theoretical assurances, is randomized smoothing truly scalable?
@@
Throughout our examination, we highlight the limitations of existing works that predominantly emphasize theoretical guarantees while overlooking the critical aspect of computational complexity, especially during inference, and dimensionality considerations. This concern holds significant weight, particularly for machine learning applications requiring minimal inference time, where randomized smoothing may fall short as a practical solution. Hence, we scrutinize each work from the perspective of scalability.
The paper is structured as follows: We begin with a comprehensive overview of the various works in the context of randomized smoothing, categorizing them into two broad areas: "Challenges of Randomized Smoothing" and "Improving Randomized Smoothing". Sections 2 and 3 delve into these categories, respectively, providing an in-depth discussion of each. Within these sections themselves we have scrutinized the works from the perspective of scalability. In the appendix we have included more detailed discussions of the works in randomized smoothing related to section 2 and section 3. In Section 4, we revisit our discussions in the preceding sections, and discuss the aspect of scalability from a real-world perspective, and provide some future directions and challenges associated with them that the community can focus as a whole.
Figure 1: This flowchart provides an overview of the main scalability challenges with RS.
2 Challenges of Randomized Smoothing
2.1 Curse of Dimensionality
@@
In their work, [8] left the case of general $l_{p}$ norms as an open problem and only discussed randomized smoothing in the context of $l_{2}$ norms. Additionally, they state that, despite the certified radius being independent of the input dimension $d$, randomized smoothing can be easily scaled to higher dimension images as in the case of higher dimension images, we can increase the variance of the noise, which will eventually lead to increase in the robust radius. Later, [4] and [22] showed that as the dimension of the input increases, it becomes increasingly difficult to defend against the general $l_{p}$ norm cases for $p>2$ as the certified radius decreases. Particularly, for the case of $l_{\infty}$, which is a common norm used in empirical attacks, the certified radius is equal to the certified radius for the $l_{2}$ norm, divided by $\sqrt{d}$. In general, there is a tight $O\left(\min\left(1,d^{\frac{1}{2}-\frac{1}{p}}\right)\right)$ dimensionality bound for $l_{p}$ norms without using any additional information other than the Neyman-Pearson technique, which cannot be improved on otherwise ([12]).
@@
We note that there has been an in-depth theoretical exploration of the issue termed the curse of dimensionality, but little work to address the issue itself. In this regards, we make note of two specific works. [47] implicitly utilise the manifold hypothesis and show that using various dimensionality reduction techniques improves robust guarantees in theory, as opposed to heuristic approaches mentioned in previous works. [34] have provided proofs and a possible solution to the issue of the curse of dimensionality using a technique called double sampling randomized smoothing (DSRS). They utilize a primary distribution in the form of a general Gaussian distribution and additional information in the form of a sampled probability from a secondary truncated distribution and, they provide a tighter radius for the $l_{2}$ and prove that their method can certify $\Theta\left(\sqrt{d}\right)$ robust radius bounds. Among these, we note that the former work can address the computational cost issue as well by working with the data in a lower dimension, but the latter introduces a computational overhead instead.
2.2 Robustness-Accuracy Trade-off
@@
In adversarial robustness literature, not just limited to RS, there is a focus on increasing the robustness bounds of a classifier. While this seems to be an optimal goal, there has been literature suggesting that increasing robustness comes with a trade-off in accuracy ([64, 56]). Both empirically and theoretically, it is seen that robustness and accuracy exist in a trade-off, contrary to what one would expect intuitively. Even in the field of RS, there have been works which have analysed the theoretical justification of the robustness-accuracy trade-off ([39, 45, 41]). Intuitively seen, randomized smoothing involves smoothing, the base classifier with Gaussian noise. While this provides certified robustness, it can be inferred that there is some loss in accuracy due to noisy inputs. Specifically, increasing the noise hyper-parameter $\sigma$ results in a larger robustly certified radius but decreased accuracy of the model. This challenge particularly limits the applicability of RS in accuracy-oriented tasks.